A hospital parking lot charges $3.00 for the first hour or part thereof, and 1.50 for each additional hour or part thereof. A weekly pass costs $42.00 and allows unlimited parking for 7 days. If each visit Johnny makes to the hospital lasts two and a half hours,what is the minimum number of visits that would make it worthwile to buy a pass?
__ visits would make it worthwile for Johnny to buy a weekly pass.

Answer by nerdybill(5408)

A hospital parking lot charges $3.00 for the first hour or part thereof, and 1.50 for each additional hour or part thereof. A weekly pass costs $42.00 and allows unlimited parking for 7 days. If each visit Johnny makes to the hospital lasts two and a half hours,what is the minimum number of visits that would make it worthwile to buy a pass?
.
If each visit is 2 and 1/2 hours:
$3 + (120/30)1.50
= 3 + 4*1.50
= 3 + 6
= $9
.
"minimum visits" = 42/9 = 4.67
Since, a "partial" visit doesn't make sense, round it up to 5.
.
Therefore:
5 visits would make it worthwile for Johnny to buy a weekly pass.